# Flow-Aware Ellipsoidal Filtration for Persistent Homology of Recurrent Signals

### v1

@@ Abstract
One common use of persistent homology is to explore the shape of point clouds, where points are assumed to be sampled from a geometric object. We propose a novel filtration, called ellipsoidal filtration, which assumes that point clouds are sampled from a dynamic smooth flow. Instead of creating topologies from point clouds at increasing scales using isotropic balls (e.g. Vietoris–Rips filtration), ellipsoidal filtration creates ellipsoids around points based on local flow variances, approximating the flow’s manifold as the scale increases. We show that constructing ellipsoidal neighbourhoods improves the denoising of recurrent signals and the estimation of recurrence times, especially when the data contain bottlenecks. Choosing ellipsoids according to the maximum persistence of the $H_{1}$ class provides a data-driven threshold for both denoising and recurrence-time estimation.
keywords:
@@
Persistent Homology , Adaptive Filters , Recurrent Signals , Ellipsoidal Filtration
[…]
[a]organization=School of Computer Science, University of Birmingham,city=Birmingham, postcode=B15 2TT, country=UK,, email= obe851@student.bham.ac.uk
[…]
[b]organization=Electrics and Electronics Department, Turkish-German University,city=Istanbul, postcode=34820, country=Turkey,, email= cihan.katar@tau.edu.tr
[…]
[c]organization=School of Computer Science, University of Birmingham,city=Birmingham, postcode=B15 2TT, country=UK,, email= m.little.1@bham.ac.uk
1 Introduction
Recurrent signals exhibit repeated patterns that manifest as close returns in their state-space trajectories, a phenomenon common in natural systems Little et al. (2007); Vieten and Weich (2020). The geometry of these trajectories encodes key information about the underlying dynamics—particularly the structure of the attractor, which indicates whether the signal is periodic, quasi-periodic, or chaotic (Kantz and Schreiber, 2003; Strogatz, 2018). In practice, the state space is reconstructed from scalar time series via time-delay (sliding-window) embeddings (Takens, 1981) or obtained directly from multivariate sensor recordings. This state space yields a discrete, high-dimensional point cloud—a set of points equipped with a distance metric.
Persistent homology, a method from applied and computational topology, characterises the topology of such point clouds across multiple spatial scales (Edelsbrunner et al., 2002; Edelsbrunner and Harer, 2008). As the scale parameter increases, data are mapped to topological spaces called simplicial complexes. The first step, a filtration, defines how points are connected at each scale. Common choices include the alpha, Čech, and Vietoris–Rips filtrations. Among these, the Vietoris–Rips filtration is often preferred for its computational and memory efficiency: points are connected when balls centred at them intersect, typically using the Euclidean distance.
@@ 1 Introduction
Time-series analysis with persistent homology largely relies on sliding-window embeddings to convert one-dimensional signals into high-dimensional point clouds. We focus on point-cloud applications in the literature; for further details readers can refer to (Ravishanker and Chen, 2019; Perea, 2019; Chazal and Michel, 2021). Perea and Harer (2015) established a theoretical link between the most persistent $H_{1}$ class, signal periodicity, and the choice of window length and embedding dimension. Gakhar and Perea (2023) further showed that sliding-window embeddings of quasi-periodic signals are dense in an $n$-torus, which can be characterised via persistent homology. Vejdemo-Johansson et al. (2015) used persistent cohomology to construct circular coordinates from sliding-window embeddings, enabling periodicity detection, dimensionality reduction, and denoising; this was later generalised to toroidal coordinates for quasi-periodic signals (Scoccola et al., 2023).
Although these methods address periodicity, they use Vietoris–Rips-based filtrations that consider only spatial relationships and treat points as samples from a static geometric object. Moreover, circular coordinates parameterise topological $S^{1}$ classes rather than directly counting cycles in time. By contrast, we adopt a close-returns perspective (Kantz and Schreiber, 2003; Little et al., 2007): replacing $n$-dimensional balls with flow-aware ellipsoids allows us to detect first returns and estimate recurrence times for each point, even when many cycles occur with rapidly varying frequency (see Section 5.4).
Fernández et al. (2023) proposed the Fermat distance as a density-aware intrinsic metric for constructing topological spaces from point clouds, which can be robust to noise. However, because it does not incorporate temporal information, we observe that it may fail to recover the main loop of the attractor in narrow bottlenecks adjacent to dense–sparse regions (Section 5.1).
Kališnik et al. (2024) presented work related in spirit to ours by introducing ellipsoidal neighbourhoods for persistent homology on static point clouds sampled from manifolds. Ellipsoids are oriented via local Principal Component Analysis (PCA) on $k$-nearest neighbours with a globally fixed aspect ratio, and simplices are added when ellipsoids intersect. Their setting is spatial rather than time-series; temporal flow and spatio–temporal structure are not modelled. Their focus is homology recovery and classification against Vietoris–Rips; by contrast, we address time-series data with a flow-aware design for denoising and recurrence-time estimation, using adaptive ellipsoids that follow the local flow variance.
Our method was developed in the search for an adaptive neighbourhood for topological denoising of quasi-periodic signals (Robinson, 2016), where prior work used $k$-NN. We explore neighbourhood choice in terms of both cardinality and geometry. The geometric neighbourhood needs to be adaptive to the local flow in recurrent dynamics. We therefore define spatio–temporal neighbourhoods and adapt ellipsoids to the local flow, evaluating denoising and recurrence-time estimation on synthetic and real accelerometer data.
@@
In our earlier work (Eryilmaz et al., 2025), we introduced an adaptive ellipsoidal filtration that incorporates local flow information by shaping anisotropic neighbourhoods according to signal gradients, enabling flow-aware persistent homology for recurrent dynamics. However, the local neighbourhood used only temporal windowing, which was insufficient to capture the variance in close returns.
@@
In the present study, we extend this framework in several ways. First, we refine the adaptive ellipsoid construction by defining spatio–temporal neighbourhoods that jointly consider local spatial proximity and short-term temporal evolution, allowing more accurate estimation of local flow geometry. Second, we apply the method to real accelerometer recordings of recurrent motion, validating its practical relevance. Third, we compare ellipsoidal filtrations with Vietoris–Rips and Fermat filtrations to evaluate differences in topological representation. Finally, we assess the utility of the resulting neighbourhoods for recurrence-time estimation, demonstrating improved performance over Vietoris–Rips-based approaches. The main contributions are:
  • We show that ellipsoidal filtration outperforms Vietoris–Rips and Fermat filtrations in capturing the main loop of the attractor under bottlenecks and varying sampling density (Section 5.1).
  • We compare neighbourhood selection strategies for topological low-pass filtering (Robinson, 2016), evaluating neighbourhood size and geometry across signal-to-noise ratios. Ellipsoidal, flow-aligned neighbourhoods reduce attenuation of low-amplitude components (Sections 5.2–5.3).
  • We revisit recurrence period density methods (Little et al., 2007): replacing isotropic balls with ellipsoids derived from the adaptive filtration, and using persistent homology to justify scale selection, improves first-return and recurrence-time estimation (Section 5.4).
@@
The paper is structured as follows. Section 2 reviews background on recurrent signals and persistent homology. Section 3 formulates the problem for temporally sampled dynamical data with observational noise and motivates flow-aware neighbourhoods. Section 4 presents the ellipsoidal filtration, our proposed method. Section 5 reports experiments on dense–sparse phase-space structure, denoising of synthetic and 3D accelerometer signals, and recurrence-time estimation. Section 6 discusses implications and limitations, and Section 7 concludes with future directions.
2 Background
2.1 Recurrent Signals
@@
$\mathbf{u}(t+\delta_{t})\in\mathbf{B}(\mathbf{u}(t),r),\quad\text{and}\quad\mathbf{u}(t+\tau)\notin\mathbf{B}(\mathbf{u}(t),r)\quad\text{for all }0<\tau<\delta_{t},$ (3)
where $\mathbf{B}(\mathbf{u}(t),r)$ is the closed ball of radius $r>0$ centred at $\mathbf{u}(t)$. The value $\delta_{t}$ is interpreted as the approximate recurrence time, indicating the system’s first return to a neighbourhood of its earlier state (Kantz and Schreiber, 2003; Little et al., 2007).
@@
This formulation naturally includes both periodic and aperiodic dynamics despite these two seems mutually exclusive. In the case of exact periodicity, recurrence times $\delta_{t}$ remain constant, and recurrence occurs at fixed intervals even as $r\to 0$. In contrast, real-world signals are generally aperiodic, and $\delta_{t}$ varies over time. In such cases, recurrence is approximate and highly dependent on the choice of neighbourhood size. The study of these variations—known as recurrence time statistics—has a long history in nonlinear time-series analysis (Kantz and Schreiber, 2003).
@@
2.2 Persistent Homology
Persistent homology is a foundational technique in topological data analysis (TDA) (Carlsson, 2009; Edelsbrunner and Harer, 2008). It identifies and tracks topological features—connected components, loops, and higher-dimensional voids—across a scale parameter $\varepsilon$.
@@ 2.2 Persistent Homology
is obtained by increasing $\varepsilon$. Exact recovery of the topology of the underlying space is often impossible with finite, noisy data, but persistent homology yields a robust multi-scale summary.
Figure 1: Illustration with a Vietoris–Rips (VR) filtration: (a) data and connections at increasing $\varepsilon$; (b) barcode showing $H_{0}$ (components) and $H_{1}$ (loops) as horizontal intervals; (c) persistence diagram (PD), mapping each feature to its birth–death point $(b,d)$. The dominant $H_{1}$ class is farthest from the diagonal.
3 Problem Statement
Point clouds are often modelled as samples from a distribution concentrated near a geometric object $X\subset\mathbb{R}^{n}$, with noise controlling the spread around $X$ (Carlsson, 2009). Under this view, common filtration techniques such as Vietoris–Rips and Čech are used to infer the Betti numbers of $X$ from a finite sample $\mathbb{X}$ (see Section 2.2). Because these constructions relate points solely through pairwise spatial distances, they impose a uniform, isotropic threshold to define neighbourhoods and build simplicial complexes. However, in point clouds derived from time-series data, the relationships between samples reflect both spatial and temporal proximity.
In this work, we consider a different setting in which the observed point cloud $\mathbb{V}\subset\mathbb{R}^{n}$ is sampled at uniform time intervals from a dynamical process with observation noise. We assume the existence of an unobserved latent trajectory $\mathbf{u}(t)$ evolving according to the same smooth flow introduced in Section 2.1, governed by Eq. (1), and we observe noisy samples given by Eq. (2).
Thus, the point cloud $\mathbb{V}=\{\mathbf{v}(t_{i})\}_{i=1}^{N}$ represents a time-parametrised dynamical process rather than a static geometric object, with its spatial density reflecting the local speed of the flow—denser in slow regions and sparser in fast ones. This anisotropic sampling invalidates the assumption that uniform spatial distances between points define meaningful neighbourhoods. Consequently, standard isotropic filtrations may fail to recover the true topological structure of the trajectory $\mathbf{u}(t)$.
In this paper, we address the following core problems, all centred on neighbourhood construction that respects the local dynamics of the flow:
  1. 1.
    designing a flow-aware filtration that approximates the manifold of the underlying trajectory using locally estimated, PCA-aligned ellipsoids instead of isotropic balls;
  2. 2.
    determining a data-driven scale, based on the death time of the dominant $H_{1}$ class, that captures the global structure while preserving local details;
  3. 3.
    performing topology-preserving denoising by averaging points within adaptive ellipsoidal neighbourhoods without distorting waveform morphology;
  4. 4.
    improving recurrence-time estimation by identifying first-return times through these adaptive, topologically consistent neighbourhoods.
4 Methods
@@
4.1 Local Covariance Estimation
The first step is to estimate how the signal varies locally around each sample. Given a point cloud representation of the signal $X=\{\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{n}\}\subset\mathbb{R}^{d}$, each point $\mathbf{x}_{i}$ represents the system’s state at a discrete time step. To capture both the temporal evolution and the recurrence structure, we construct two types of local neighbourhoods:
  1. 1.
    an index-based (temporal) neighbourhood, which captures short-term evolution in time, and
  2. 2.
    a state-space (spatial) neighbourhood, which identifies recurrent states that are close in phase space but possibly far apart in time.

### v2

@@ Abstract
Recurrent signals give rise to trajectories that repeatedly return close to earlier states in state space. Many analysis methods therefore require a principled notion of similarity between states. In practice, a recurrence threshold sets the scale of the neighbourhood used to define when two states are considered close. Close returns can also support topology-preserving denoising in state space, aiming to reduce noise while preserving the trajectory’s structure, which classical denoising methods may distort. The effectiveness of both denoising and recurrence analysis therefore depends critically on how these neighbourhoods are modelled and scaled.
This work introduces a flow-aware ellipsoidal filtration for persistent homology based on a spatio–temporal covariance construction that estimates local flow geometry from both temporal and spatial neighbours. Unlike isotropic constructions based on balls (e.g. the Vietoris–Rips filtration), the proposed method assigns an ellipsoid to each point, with orientation and axis lengths determined by local flow variances. When a dominant $H_{1}$ feature reflects the recurrent loop structure, its persistence interval provides a data-driven scale selection. Across the considered experiments, flow-aware ellipsoidal neighbourhoods improve topology-preserving denoising and first-recurrence-time estimation relative to the Vietoris–Rips filtration. Overall, the results indicate that persistent homology can be more informative for dynamical systems when domain knowledge is used to incorporate anisotropy.
@@
Recurrent signals; recurrence threshold; persistent homology; flow-aware ellipsoidal filtration; topology-preserving denoising
preprint: AIP/123-QED
Persistent homology is increasingly used to analyse time-series data from dynamical systems. However, most applications rely on isotropic distance-based neighbourhoods, which may not capture the local flow structure of a trajectory. In this paper, we construct flow-aware ellipsoidal neighbourhoods for each data point using local spatio–temporal covariance. Persistent homology is then used to guide the scaling of these neighbourhoods, which are applied to topology-preserving denoising and close-return detection. The resulting neighbourhoods more faithfully preserve trajectory geometry and improve recurrence-time estimation in the presence of narrow bottlenecks.
I Introduction
Recurrent behaviour is common in nonlinear systems and forms the basis of recurrence-based analysis methods (Webber and Marwan, 2015; Marwan and Kraemer, 2023). To use these returns in practice, one must specify when two states should be considered close, which is typically done by choosing a recurrence threshold that sets the neighbourhood scale for proximity (Schinkel et al., 2008; Zou et al., 2019; Li et al., 2025). This choice is important because the same neighbourhood definition can be used both to quantify recurrence structure (Webber and Marwan, 2015) and to denoise in state space by local averaging (Robinson, 2016), while aiming to preserve the trajectory’s overall shape. By contrast, standard denoising methods may distort this structure when applied without reference to state-space geometry. For these reasons, both recurrence analysis and topology-preserving denoising depend critically on how neighbourhoods are defined and how their scale is selected from data.
Persistent homology (Edelsbrunner et al., 2002; Edelsbrunner and Harer, 2008; Carlsson and Vejdemo-Johansson, 2021), a tool from applied and computational topology, provides a multiscale description of the geometric and topological structure of these sampled trajectories. Instead of analysing the data at a single resolution, it records how topological features emerge and disappear as a neighbourhood scale parameter increases. This process is governed by a filtration, which specifies how local neighbourhoods around data points grow and intersect across scales. Common constructions include the alpha, Čech, and Vietoris–Rips filtrations. In practice, the Vietoris–Rips filtration is frequently used because of its computational simplicity, defining connectivity through intersections of isotropic balls under a chosen distance metric.
@@ I Introduction
Time-series analysis with persistent homology largely relies on sliding-window embeddings to convert one-dimensional signals into high-dimensional point clouds. We focus on time-series analysis via point-cloud representations; readers seeking further examples of TDA applications may consult the DONUT repository (Giunti et al., 2022). Perea and Harer Perea and Harer (2015) established a theoretical link between the most persistent $H_{1}$ class, signal periodicity, and the choice of window length and embedding dimension. Gakhar and Perea Gakhar and Perea (2024) further showed that sliding-window embeddings of quasi-periodic signals are dense in an $n$-torus, which can be characterised via persistent homology. Vejdemo-Johansson Vejdemo-Johansson et al. (2015) used persistent cohomology to construct circular coordinates from sliding-window embeddings, enabling periodicity detection, dimensionality reduction, and denoising; this was later generalised to toroidal coordinates for quasi-periodic signals (Scoccola et al., 2023).
More broadly, persistent homology has been applied to dynamical-systems data to study delay-coordinate reconstructions of chaotic attractors (Maletić et al., 2016), to guide embedding-delay selection (SToPS) (Tan et al., 2023), and to assess whether learned models reproduce reference dynamics from sampled trajectories (Tan et al., 2021). Complementing these point-cloud-based approaches, cubical homology on phase-portrait images yields topological descriptors (e.g. Betti curves and persistence-based features) for automated analysis and classification of dynamical behaviours (Hussain Shah et al., 2025). Most of the above constructions use Vietoris–Rips filtrations under an ambient Euclidean metric; under the manifold assumption, this motivates alternatives that modify the neighbourhood relation, such as density-aware intrinsic metrics (Fermat distance) (Fernández et al., 2023) or temporally informed anisotropic neighbourhoods.
A further motivation for anisotropic neighbourhoods comes from the manifold hypothesis, which posits that high-dimensional data may lie near a lower-dimensional manifold and motivates algorithmic tests of such structure (Fefferman et al., 2016). In this setting, isotropic balls can be a crude local model, since the data may vary primarily along tangent directions (Marchetti, 2025). Breiding et al. proposed replacing $\varepsilon$-balls by ellipsoids aligned with estimated tangent spaces (obtained, when polynomial equations are available, from Jacobian information), and reported longer persistence bars in their examples (Breiding et al., 2018). Complementing this, Kališnik and Lešnik proved that unions of tangent-elongated ellipsoids can deformation retract to a manifold under suitable sampling assumptions, with improved sample-density requirements over ball thickenings (Kališnik and Lešnik, 2024). Canova et al. Canova et al. (2025) define an ellipsoid complex for point clouds sampled from manifolds. Each point is assigned an ellipsoid whose orientation is given by local principal component analysis (PCA) on $k$-nearest neighbours and whose axis ratios are fixed globally; simplices are added when the corresponding ellipsoids intersect. These ellipsoidal constructions were developed for manifold reconstruction or shape classification, where persistence diagrams are converted into feature vectors for classification tasks. Neighbourhoods are defined using purely spatial criteria, without incorporating temporal ordering of the samples.
@@
While these methods operate on static point clouds, our work addresses point clouds sampled from dynamical systems. It is motivated by the search for adaptive, data-driven neighbourhoods for topological denoising of quasi-periodic signals (Robinson, 2016), rather than relying on fixed $k$-nearest-neighbour ($k$-NN) constructions. In contrast to these spatial constructions, in our earlier work (Eryilmaz et al., 2025), we introduced an adaptive ellipsoidal filtration that incorporates local flow information by shaping anisotropic neighbourhoods according to signal gradients, thereby enabling flow-aware persistent homology for recurrent dynamics. However, in that formulation the local neighbourhood was defined purely by temporal windowing, which limited its ability to capture variability in close returns.
@@
In the present study, this framework is extended in several ways. First, the adaptive ellipsoid construction is refined by defining spatio–temporal neighbourhoods that jointly consider local spatial proximity and short-term temporal evolution, allowing more accurate estimation of local flow geometry. Second, the method is applied to real accelerometer recordings of recurrent motion, validating its practical relevance. Third, ellipsoidal filtrations are compared with Vietoris–Rips and Fermat filtrations to evaluate differences in topological representation. Finally, the utility of the resulting neighbourhoods for recurrence-time estimation is assessed, and improved performance over Vietoris–Rips-based approaches is demonstrated. The main contributions are:
  • A flow-aware geometric interpretation of point-cloud density for recurrent dynamics is introduced. For uniformly sampled trajectories, sparse regions of the point cloud correspond to fast-flow segments, and dense regions to slow-flow segments. Around each point, a flow-aware ellipsoidal neighbourhood region is attached, whose size and orientation reflect the local flow, and points are connected when their ellipsoids intersect. In the experiments considered here, the ellipsoidal filtration recovers the main loop in attractors with bottlenecks (i.e. narrow pinch regions of the loop), whereas Vietoris–Rips and Fermat filtrations tend to split this into two loops (Section V.1).
  • It is shown that the ellipsoidal filtration provides a principled and interpretable guideline for neighbourhood selection in topological low-pass filtering of recurrent signals (Robinson, 2016). We assume a suitable embedding in which recurrent behaviour manifests as loop structures in state space, so that significant $H_{1}$ features correspond to this recurrence. The death scale of the most persistent $H_{1}$ class is used as a conservative upper bound for neighbourhood selection, ensuring that the dominant loop structure remains observable. Empirically, we find that the variance-adaptive nature of the flow-aware ellipsoidal neighbourhoods can preserve the underlying geometric structure even slightly beyond this bound. Compared with $k$-NN and Vietoris–Rips-based isotropic neighbourhoods, the proposed flow-aware ellipsoids more consistently preserve fine-scale structures across bottlenecks (Sections V.2V.3).
  • Recurrence period density methods (Little et al., 2007) are revisited, and a topologically grounded way to choose recurrence threshold is provided. Vietoris–Rips-based isotropic balls are compared with flow-aware ellipsoids, with both scaled using guidance from persistent homology, specifically the scale of the dominant $H_{1}$ class. The resulting flow-aware ellipsoidal recurrence neighbourhoods improve first-return time estimation compared with Vietoris–Rips-based constructions (Section V.4).
@@
The paper is structured as follows. Section II reviews background on recurrent signals and persistent homology. Section III formulates the problem for temporally sampled dynamical data with observational noise and motivates flow-aware neighbourhoods. Section IV presents the construction of the flow-aware ellipsoidal filtration. Section V reports experiments on dense–sparse phase-space structure, denoising of synthetic and 3D accelerometer signals, and recurrence-time estimation. Section VI discusses implications, limitations and future directions, and Section VII concludes.
II Background
This section introduces recurrence in state space and summarises persistent homology, establishing the notation used throughout the paper.
II.1 Recurrent Signals
@@
[…]
where $\mathbf{B}(\mathbf{u}(t),r)$ is the closed ball of radius $r>0$ centred at $\mathbf{u}(t)$. The value $\delta_{t}$ is interpreted as the approximate recurrence time, indicating the system’s first return to a neighbourhood of its earlier state (Kantz and Schreiber, 2003).
@@
This formulation naturally includes both periodic and aperiodic dynamics despite these two seeming mutually exclusive. In the case of exact periodicity, recurrence times $\delta_{t}$ remain constant, and recurrence occurs at fixed intervals even as $r\to 0$. In contrast, real-world signals are generally aperiodic, and $\delta_{t}$ varies over time. In such cases, recurrence is approximate and highly dependent on the choice of neighbourhood size. The study of these variations—known as recurrence time statistics—has a long history in nonlinear time-series analysis (Kantz and Schreiber, 2003).
@@
II.2 Persistent Homology
Persistent homology is a foundational technique in topological data analysis (TDA) (Carlsson and Vejdemo-Johansson, 2021; Chazal and Michel, 2021). It identifies and tracks topological features—connected components, loops, and higher-dimensional voids—across a scale parameter $\varepsilon$.
@@ II.2 Persistent Homology
is obtained by increasing $\varepsilon$. Exact recovery of the topology of the underlying space is often impossible with finite, noisy data, but persistent homology yields a robust multi-scale summary; this procedure is summarised in Fig. 1.
Figure 1: Illustration with a Vietoris–Rips (VR) filtration: (a) data and connections at increasing $\varepsilon$; (b) barcode showing $H_{0}$ (components) and $H_{1}$ (loops) as horizontal intervals; (c) persistence diagram (PD), mapping each feature to its birth–death point $(b,c)$. The dominant $H_{1}$ class is farthest from the diagonal.
III Problem Statement
Point clouds are often modelled as samples from a distribution concentrated near a geometric object $X\subset\mathbb{R}^{n}$, with noise controlling the spread around $X$ (Carlsson, 2009). Under this view, common filtration techniques such as Vietoris–Rips and Čech are used to infer the Betti numbers of $X$ from a finite sample $\mathbb{X}$ (see Section II.2). Because these constructions relate points solely through pairwise spatial distances, they impose a uniform, isotropic threshold to define neighbourhoods and build simplicial complexes. However, in point clouds derived from time-series data, the relationships between samples reflect both spatial and temporal proximity.
In this work, a different setting is considered in which the observed point cloud $\mathbb{V}\subset\mathbb{R}^{n}$ is sampled at uniform time intervals from a dynamical process with observation noise. It is assumed that there exists an unobserved latent trajectory $\mathbf{u}(t)$ evolving according to the smooth flow introduced in Section II.1, governed by Eq. (1), and that noisy samples are observed according to Eq. (2).
Thus, the point cloud $\mathbb{V}=\{\mathbf{v}(t_{i})\}_{i=1}^{N}$ represents a time-parametrised dynamical process rather than a static geometric object, with its spatial density reflecting the local speed of the flow—denser in slow regions and sparser in fast ones. This anisotropic induced spatial density invalidates the assumption that uniform spatial distances between points define meaningful neighbourhoods, and standard isotropic filtrations may therefore fail to recover the true topological structure of the trajectory $\mathbf{u}(t)$.
The core problem addressed in this paper is to construct neighbourhoods and filtrations that respect the local dynamics of the flow. Specifically, the aim is to approximate the manifold of the underlying trajectory using locally estimated, PCA-aligned ellipsoids instead of isotropic balls; to select a data-driven scale, based on the dominant $H_{1}$ class, that captures global structure while preserving local detail; and to exploit these flow-aware neighbourhoods both for topology-preserving denoising and for improved recurrence-time estimation.
IV Methods
@@
IV.1 Local Covariance Estimation
The first step is to estimate how the signal varies locally around each sample. Given a point cloud representation of the signal $X=\{\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{n}\}\subset\mathbb{R}^{d}$, each point $\mathbf{x}_{i}$ represents the system’s state at a discrete time step. To capture both the temporal evolution and the recurrence structure, two types of local neighbourhoods are used: (i) an index-based (temporal) neighbourhood, which captures short-term evolution in time, and (ii) a state-space (spatial) neighbourhood, which identifies recurrent states that are close in state-space but possibly far apart in time.